### v4

# Database for the meta-analysis of the social cost of carbon (v2026.1)

Abstract
A new version of the database for the meta-analysis of estimates of the social cost of carbon is presented. New records were added.
Keywords: meta-analysis, social cost of carbon
JEL codes: Q54, Y10, Z13
1 Introduction
This paper describes the databases underlying the meta-analysis of the social cost of carbon.
Previous meta-analyses of the social cost of carbon were published by Tol (2005b), van den Bergh and Botzen (2014), Havránek et al. (2015), Wang et al. (2019) and Moore et al. (2024). My first attempt was updated in Tol (2008, 2009, 2010b, 2011, 2013b, 2014, 2015, 2018, 2019a, 2020, 2023a, 2023b), Anthoff and Tol (2022), Tol (2025b) and Tol (2025a). Many of the updates involve newer data only. The exceptions are Tol (2008) and Anthoff and Tol (2022), who focus on the tail of the distribution, Tol (2018) and Tol (2025a) who try to detect publication bias (the latter successfully), and Tol (2008) and Tol (2023b), who study trends. Tol and Tol (2023) is about software for online visualization of meta-analysis.
The new version of the database includes new records.
Section 2 presents the databases. Section 3 shows descriptive statistics. Section 4 concludes.
2 Data
2.1 Literature
The current version of the database includes 528 papers, published in 156 journals, spanning the period from 1980 to 2025. These papers were collected over two decades. Only two papers are missing, the others are in the author’s library, mostly in electronic form some only in hard copy. Papers were collected by systematic literature search, personal acquaintance, and exploration of the citations of previously considered papers.
This version of the database adds 60 papers published in 2025: Akdağ (2025), Aksoy and Bayram Arlı (2025), Arab et al. (2025), Ashenfarb and Kotchen (2025), Behzadi and Arani (2025), Bhattacharya et al. (2025), Bhave et al. (2024), Brooks et al. (2025), Caesary et al. (2025), Castellini et al. (2025), Chapin et al. (2025), Chen et al. (2025a, b), Cheu and Gammans (2025), Cho et al. (2025), Coppens et al. (2025), Cunanan et al. (2025), da Costa et al. (2025), Deng et al. (2025), Eduardo et al. (2025), Estrada et al. (2025a, b), Fell et al. (2025), Feng (2025), Folini et al. (2025), García et al. (2025), Islam et al. (2025), Kashanian et al. (2025), Khan et al. (2025), Kerns et al. (2025), Kotchen (2025), Laamouri and Khattabi (2025), Landry (2025), Lima et al. (2025), Lubis et al. (2025), Manan et al. (2025), Mohammadyari et al. (2025), Naef et al. (2025), Neal et al. (2025), Ovaere and Proost (2025), Piazza et al. (2025), Van Der Ploeg (2025), Pranita and Sarjana (2025), Rao et al. (2025), Rose et al. (2025), Schaumann and de Asenjo (2025), Serfas (2025), Springmann et al. (2025), Sreelekshmi et al. (2025), Wang et al. (2025), Wu (2025), Wu et al. (2025a, b), Xiao et al. (2025), Xu et al. (2025), Xuan et al. (2025), Yu et al. (2025), Zhang et al. (2025), Zhu et al. (2025).
I also added 21 previously overlooked papers: Al-Amin et al. (2015), Aleshina et al. (2024), Bilal and Känzig (2024), Government of Canada (2022), Deng et al. (2024), Sylvan (2017b), Eteriki et al. (2023), GAO (2020), Howard and Sylvan (2015a, b), Landry (2021), Li et al. (2025), Rasiah et al. (2016), Raynaud (2014), Rose et al. (2022), Sarkar et al. (2019), Shelanski and Obstfeld (2015), Siagian et al. (2024), Sylvan (2017a), Umweltbundesamt (2024), Wong et al. (2016).
These paper add to 446 papers included in the previous version of the meta-analysis: Ackerman and Finlayson (2006), Ackerman and Munitz (2012, 2016), Ackerman and Stanton (2012), Ackerman et al. (2013), Adler et al. (2017), Agliardi and Xepapadeas (2022), Allen (2016), Anderson et al. (2014), Anthoff (2007, 2009), Anthoff and Emmerling (2019), Anthoff and Tol (2009, 2010, 2011, 2013, 2014, 2022), Anthoff et al. (2009a, b, c, 2011a, 2011b), Asplund (2017, 2019), Audus (1998), Ayres and Walter (1991), Azar (1994), Azar et al. (2023), Azar and Sterner (1996), Balmford et al. (2023), Barnett et al. (2020), Bastien-Olvera and Moore (2021), Barrage (2014, 2018, 2020b, 2020a), Barrage and Nordhaus (2023, 2024), Bauer and Rudebusch (2023), Belfiori and Rezai (2024), von Below (2014), Berger et al. (2017), Bertram et al. (2024), Bherwani et al. (2019, 2024), van den Bijgaart et al. (2013, 2016), Bond (2024), Botzen and van den Bergh (2012), Brander et al. (2010), Braun et al. (2024), Brausmann and Bretschger (2024), van den Bremer and van der Ploeg (2021), Bressler (2021), Bretschger and Karydas (2018), Bretschger and Pattakou (2019), Brock and Xepapadeas (2017, 2019), Budolfson and Dennig (2020), Budolfson et al. (2017, 2019), Caesary et al. (2023), Cai and Lontzek (2019), Cai et al. (2012, 2015, 2016), Calvas et al. (2024), Ceronsky et al. (2005, 2011), Chung et al. (2024), Clarkson and Deyes (2002), Cline (1992, 1997, 2004), Coleman et al. (2021), Coppola et al. (2024), Crost and Traeger (2013, 2014), Czyz and Safarzyńska (2023), Dangl and Wirl (2007), Daniel et al. (2019), Dayaratna and McKitrick (2023), Dayaratna et al. (2017, 2020), Dennig (2014), Dennig et al. (2015), Dietz (2011), Dietz and Koninx (2022), Dietz et al. (2021a, b), Dietz and Stern (2015), Dietz and Venmans (2019), Dong et al. (2024), Downing et al. (1996, 2005), Drupp and Hänsel (2021), Ekholm (2018), EPA (2023), EPA and NHTSA (2009), Espagne et al. (2018), Eyre et al. (1999), Fankhauser and Pearce (1993), Fankhauser (1994, 1995), Farooq and Cora (2024), Faulwasser et al. (2018b, a), Fiestas-Chevez et al. (2024), Fillon et al. (2023), Foley et al. (2013), Freeman and Groom (2015), Freeman et al. (2015), Freeman and Groom (2016), Fu et al. (2023), Gerlagh (2023), Gerlagh et al. (2020), Gerlagh and Liski (2017, 2018), Gerlagh et al. (2023), Gillingham et al. (2018), Glotter et al. (2014), Golosov et al. (2014), Golub and Brody (2017), González-Eguino and Neumann (2016), Gössling and Humpe (2024), Goulder and Mathai (2000), Greenstone et al. (2013), Griffiths et al. (2012), Groom et al. (2007), Groom and Venmans (2023), Groshans et al. (2019), Gschnaller (2020), Guivarch and Pottier (2018), Guo et al. (2006), Hafeez et al. (2017), Haites (1998), Hambel et al. (2021b, a, 2024), Hanley and Tinch (2004), Hänsel and Quaas (2018), Hänsel et al. (2020, 2021), Haraden (1992, 1993), Hasan et al. (2024), Hassler et al. (2016), Hastunç et al. (2025), Hatase and Managi (2015), Heal and Millner (2014), Heinzow and Tol (2003), Hepburn et al. (2009), Hillebrand and Hillebrand (2019, 2023), Hoffman et al. (2024), Hohmeyer (1996, 2002, 2004), Hohmeyer and Gaertner (1992), Hong et al. (2023), Hope (2005a, b, c, 2006a, 2006b, 2006c, 2008a, 2008b, 2009, 2011b, 2011a, 2013a, 2013b, 2020), Hope and Hope (2013), Hope and Maul (1996), Howard and Sterner (2017), Howard and Sylvan (2020), Howarth (1998), Howarth et al. (2014), Hwang et al. (2013, 2016, 2017), Hwang (2017), Hwang et al. (2019), Interagency Working Group on the Social Cost of Carbon (2010, 2013, 2015, 2021), Ibrahim et al. (2024), Ikefuji et al. (2021), Iverson (2012), Iverson et al. (2015), Iverson and Karp (2021), Jaakkola and Millner (2022), Jaakkola and van der Ploeg (2019), Hoffman et al. (2024), Al-Jabir and Isaifan (2024), Jensen and Traeger (2013, 2014, 2021), Jiang et al. (2024), Jin et al. (2024), Johnson and Hope (2012), Kalkuhl and Wenz (2020), Karydas and Xepapadeas (2019), Kaushal and Navrud (2023), Kellett et al. (2019), Khabarov et al. (2022), Kelly and Kolstad (2001), Kemfert and Schill (2010), Kessler (2017), Kikstra et al. (2021), Knoke et al. (2023), Koch and Leimbach (2023), Kon and Caner (2024), Kopp et al. (2012), Kotchen (2018), Kotlikoff et al. (2021), Krewitt and Schlomann (2006), Kulkarni et al. (2024), Lemoine (2015, 2021), Lemoine and Traeger (2014, 2016a, 2016b), Li et al. (2016), Link and Tol (2004), Lintunen and Rautiainen (2021), Liu et al. (2022), Lontzek et al. (2015), Loube (2024), Lucchesi et al. (2024), Lupi and Marsiglio (2021), Maddison (1994, 1995), Manne (2004), Mardones (2024), Marten (2011, 2014), Marten et al. (2015), Marten and Newbold (2012a, b, 2013), Mastandrea and Schneider (2001), Mendelsohn (2003, 2004, 2005), Meng et al. (2024), van der Mensbrugghe (2023), Mikhailova et al. (2019, 2024b, 2024a), MnPUC (2018), Molocchi and Mela (2024), Moore and Diaz (2015), Moore et al. (2017), Moyer et al. (2014), Muangjai et al. (2024), Naeini et al. (2020), Narita et al. (2009, 2010a, 2010b), Newbold et al. (2010, 2013), Newbold and Marten (2014), Newell and Pizer (2003, 2004), Newell et al. (2022), Niu and Zou (2024), Nordhaus (1980, 1982, 1989, 1990a, 1990b, 1991a, 1991b, 1991c, 1992, 1993a, 1993b, 1993c, 1994b, 1994a, 1997, 2001, 2007b, 2007c, 2007a, 2008, 2009, 2010, 2011, 2013a, 2013b, 2014, 2015, 2017b, 2017a, 2018b, 2018a, 2019b, 2019a), Nordhaus and Boyer (2000), Nordhaus and Popp (1997), Nordhaus and Sztorc (2013), Nordhaus and Yang (1996), NYSERDA and RFF (2021), Okullo (2020), Olijslagers (2022), Olijslagers et al. (2023), Olijslagers and van Wijnbergen (2024), Ortiz et al. (2011), Otto et al. (2013), Parry (1993), Pearce (2003, 2005), Peck and Teisberg (1992, 1993, 1994, 1995, 1996), Penner et al. (1992), Perrissin Fabert et al. (2012), Pezzey and Burke (2014), Pindyck (2017, 2019), Piontek et al. (2019), Pizer (1999, 2002), Plambeck and Hope (1996), van der Ploeg (2014, 2015), der Ploeg and Rezai (2016), van der Ploeg and Rezai (2017), van der Ploeg and Rezai (2018), van der Ploeg and Rezai (2019b, a), van der Ploeg et al. (2012), van der Ploeg (2018), van der Ploeg and Rezai (2021), van der Ploeg (2021), van der Ploeg and de Zeeuw (2013), de Zeeuw and van der Ploeg (2014), van der Ploeg and de Zeeuw (2016, 2017, 2019), Poelhekke (2019), Popp (2004), Portland and South Portland (2021), Pottier et al. (2015), Prest and Stock (2023), Prest et al. (2024), Price et al. (2007), Pulhin et al. (2024), Pycroft et al. (2011, 2014), Quiggin (2018), Rautiainen and Lintunen (2017), Reilly and Richards (1993), Rennert and Kingdon (2019), Rennert et al. (2021, 2022), Rezai (2010), Rezai and van der Ploeg (2015, 2016, 2017b, 2017c, 2017a), Rezai et al. (2012, 2020), Ricke et al. (2018), Rickels et al. (2023, 2024), Rose et al. (2017), Roughgarden and Schneider (1999), Rudik (2020), Russell et al. (2022), Di Russo et al. (2025), Safarzyńska and van den Bergh (2022), Schauer (1995), Schmidt et al. (2013), Schultes et al. (2021), Scovronick et al. (2017, 2020), Shafiei-Alavijeh et al. (2024), Sen et al. (2024), Shindell (2015), Shindell et al. (2017), Smith et al. (2023), Sohn (2019), Sohngen (2010), Stern et al. (2006), Stern and Taylor (2007), Striepe et al. (2024), Su et al. (2017, 2024), Taconet et al. (2021), Tao et al. (2024), Tian et al. (2019), Tibebu et al. (2021, 2024), Tol (1999, 2003, 2005a, 2005c, 2010a, 2010c, 2012b, 2013a, 2013c, 2019b), Tol and Downing (2000, 2001), Tol (2020, 2024), Traeger (2014, 2015, 2023), Umweltbundesamt (2007, 2012, 2019, 2020), Uzawa (2003), Wahba and Hope (2006), Waldhoff et al. (2011, 2014), Wang et al. (2022, 2024), Watkiss (2005), Watkiss and Downing (2008), Weitzman (2013), Weller et al. (2015), van der Wijst et al. (2021), van der Wijst et al. (2023), Wong et al. (2015), Wu et al. (2024), Yang et al. (2018), Yilma and Yitay (2024), Yohe and Hope (2013), Yoo and Mendelsohn (2018, 2021), Zhao et al. (2023), Zhen and Tian (2020) and Zhen et al. (2018).
2.2 Main database
Results and assumptions were coded for each of the 528 papers. In most cases, tables and text were used. Some papers present results only graphically, in which case authors were contacted by email. If there was no response—in one case, the authors responded that their hard disk had crashed—graphical results were digitized with the grabit function in Matlab.
The main database is in Excel. It consists of the following fields, for each of the 14,884 estimates of the social cost of carbon:
  • 1.
    Year of publication.
  • 2.
    Author weight, indicating the apparent weight the authors give to each estimate.
  • 3.
    Paper weight, equal to one over the number of estimates contained in the paper.
  • 4.
    Quality weight, equal to one plus:
    • (a)
      Peer-review, a dummy variable indicating whether the paper was independently reviewed;
    • (b)
      Method, a dummy variable indicating whether the social cost of carbon was calculated or approximated using a mathematically accepted method; and
    • (c)
      Scenario, a dummy variable indicating whether not-implausible scenarios were used.
  • 5.
    Censor, a variable equal to one if the estimate is lower than \$1,594/tC, the Leviathan tax in 2019 (Tol, 2012a), and zero if it is greater than \$11,571/tC, the maximum ability to pay. The variable is scaled linearly between 0 and 1 for values in between.
  • 6.
    Social cost of carbon, as reported.
  • 7.
    Emission year.
  • 8.
    Dollar year.
  • 9.
    Unit, indicating whether the social cost of carbon is expressed in metric tonnes [^1] No study reports non-metric units. of carbon or of carbon dioxide.
  • 10.
    Legacy The social cost of carbon for 2010 in 2010 US dollar per metric tonne of carbon.
  • 11.
    The social cost of carbon for 2025 in 2024 US dollar per metric tonne of carbon.
  • 12.
    Negative, a dummy variable whether or not the study permits the possibility of a social benefit of carbon.
  • 13.
    Consumption discount rate.
  • 14.
    Pure rate of time preference.
  • 15.
    Rate of risk aversion.
  • 16.
    Inverse of the elasticity of intertemporal substitution.
  • 17.
    Rate in inequity aversion.
  • 18.
    Assumed level impact of 2.5 warming.
  • 19.
    Functional form of the level damage function.
  • 20.
    Assumed growth impact of 1.0 warming.
  • 21.
    Functional form of the growth damage function.
  • 22.
    Global warming in 2100.
  • 23.
    Equity weights, a dummy variable whether or not used.
  • 24.
    Mean, a dummy variable whether the expectation of the social cost of carbon is reported (1) or some other statistic (0).
  • 25.
    Stochastic, a dummy variable whether the model used is stochastic (1) or not.
  • 26.
    Analytic, a dummy variable whether the model is solved analytically (1) or numerically (0).
  • 27.
    Dynamic, a variable indicating whether vulnerability to climate change falls with development (1), increases (-1), or does not change (0).
  • 28.
    Pigou, a dummy variable whether the social cost of carbon is imposed on emissions or not.
  • 29.
    Paper ID.
The data are on GitHub, file socialcostcarbon.xlsx.
A supplementary database records the social cost of carbon for different years of emission. This is used to calculate the population average growth rate, which is 2.07% per year in the current database. This growth rate is used to shift the social cost of carbon for the year of reporting to 2025.
A study-specific growth rate is computed too. This is used, where available, to compute an alternative social cost of carbon for 2025.
These data are in the same file as above.
2.3 Satellite data
There are four satellite databases:
  1. 1.
    Bibliography, in BibTex. This preserves the sanity of the author and allows others to replicate the research. This database contains:
    • (a)
      Paper ID.
    • (b)
      Title.
    • (c)
      Authors.
    • (d)
      Publication year.
    • (e)
      Journal name / volume title / publication series.
    The data are on GitHub, file scc.bib
  2. 2.
    Co-author network, in Matlab and Excel. This allows for testing idiosyncratic preferences and influence. This database contains:
    • (a)
      Paper ID.
    • (b)
      Author ID.
    • (c)
      Gender (at the time of publication).
    • (d)
      A dummy variable indicating whether a particular researcher authored a particular paper.
    The data are on GitHub, files citation.m and connect.xlsx.
  3. 3.
    Country of affiliation, in Excel. This allows for testing representativeness. This database contains:
    • (a)
      Paper ID.
    • (b)
      Country ID.
    • (c)
      A dummy variable indicating whether a particular paper was authored by a researcher based in a particular country.
    The data are on GitHub, file connect.xlsx
  4. 4.
    Citation network, in Matlab. This allows for testing influence and citation bias. This database contains:
    • (a)
      Paper ID.
    • (b)
      Citation ID.
    • (c)
      Citation ID dyads.
    The data are on GitHub, file citation.m
The five databases have the field Paper ID in common and can therefore readily be combined in any constellation.
3 Descriptive statistics
Figure 1 shows the number of papers and the number of estimates per publication year. The first two papers were published in 1980 and 1982, with 12 estimates between them. More papers were published in the early 1990s. Since 2010, ten papers or more were published per year. The record was set in 2025 with 62 new papers. The number of estimates increased more rapidly than the number of papers as computers got faster, referees more demanding, and supplementary information more extensive. The maximum was reached in 2021, with 5,511 estimates reported (in 27 papers). [^2] There are many more unpublished results. Chris W. Hope routinely runs Monte Carlo analysis with the page model but never reports the full results. Rudik (2020) is based on 50,000 Monte Carlo runs, Anthoff and Tol (2022) on 100,000 each for three alternative models.
Figure 2 shows the histogram of all estimates of the social cost of carbon. The mode of the empirical distribution lies between \$50/tC and \$75/tC. A second mode lies between \$900/tC and \$925/tC, the oft-repeated estimate of Rennert et al. (2022). The distribution is skewed, with a pronounced right tail. The mean is thus much larger than the mode: \$411/tC (s.e. \$10/tC). The number of estimates showing a social benefit of carbon is small, only 1.1% of the total.
Figure 3 shows the mean and range of estimates of the social cost of carbon over time. Early estimates were very high. Later estimates were lower but quite variable from year to year. There appears to be an upward trend from 2010 to 2023.
A refined field reveals that 81.5% of estimates of the social cost of carbon assume that vulnerability to climate change is independent of development (cf. Schelling, 1984). 17% (1.5%) assume that vulnerability falls (rises) with income. The social cost of carbon is \$124/tC lower (\$47/tC higher) than with constant vulnerability.
Eight hundred seventy-four authors contributed to the literature on the social cost of carbon. The number of authors per paper varies between 1 and 29 (Bertram et al., 2024). The average paper has 1.66 authors. Figure 4 shows the 15 authors who contributed most to the literature on the social cost of carbon. Papers are shared equally between authors. That is, if a paper has two (three) authors, both (each) get assigned a half (third) paper, so that the sum of author-contributions equals the number of papers. Richard S.J. Tol (co-)authored most papers (46) but William D. Nordhaus, who tends to work alone or with one junior, has made the largest contribution (6.8%). The ten most prolific authors together cover only 25.5% of the literature.
The average estimate for female authors is \$51/tC higher than for males. The average estimate in economics journals is \$137/tC lower than in other journals.
Figure 5 shows the same information but for the country of affiliation. Some 38.5% of all papers on the social cost of carbon were authored by researchers based in the USA, 17.3% in the UK, 7.8% in Germany, and 5.0% in China. Forty-five other countries contributed to this literature. The remaining 140-odd countries did not. There are few papers from authors in Latin America and Africa.
Figure 7 shows the co-author networks of everyone who has published five papers or more on the social cost of carbon. Authors of six papers or more are named. These seven networks together include N out of 872 authors. The smallest network is Stephen C. Peck and Thomas J. Teisberg. The largest network has N members, including most of the most prolific authors (cf. Figure 4).
Figure 8 shows the citation network. The node size is weighted arithmetic incloseness. Incloseness is a measure of influence, counting not just the citations of a paper, but also citations of citations, citations of citations of citations, and so on. Incloseness is an average distance. In this case, the arithmetic average is used, or rather a weighted average where the weights equal one over the number of citations. That is, if a paper is the only citation in another paper, the weight is one. If a paper is one of two citations, the weight is one-half. And so on. Figure 8 reveals a dense and complicated citation network. Because papers are often circulated as working papers prior to publication, there are cycles, short ones—a paper citing a paper that cites it—as well as long ones—up to nine papers. The most influential papers are named. These were identified by running a regression of incloseness on the year of publication. The 5% papers that are furthest from the regression line are named.
4 Conclusion
This paper presents the 2026 version of the database of estimates of the social cost of carbon. Some records were updated and 81 new records added. The 2026 database is considerably larger than the 2025 one: 528 vs 446 papers. The average social cost of carbon of previously included studies is \$400/tC; the new studies average \$629/tC.
The database will be updated with newer publications on the social cost of carbon and inadvertently overlooked older publications. New fields may be added as necessary.
Figure 1: The number of papers on and estimates of the social cost of carbon by year.
Figure 2: The histogram of published estimates of the social cost of carbon. Estimates are author- and quality-weighted and censored.
Figure 3: The average social cost of carbon by year of publication. The interval shown is the mean plus and minus the standard deviation. Estimates are author- and quality-weighted and censored. Both panels show the same data, but the right panel uses an abridged vertical axis.
Figure 4: Contribution to the literature on the social cost of carbon by author.
Figure 5: The number of papers on the social cost of carbon by country of affiliation.